OpenStudy (emilyjones284):

Help please ill give a medal (:

OpenStudy (emilyjones284):

What is the solution of the equation \[4^{x+1}=21\]?

hartnn (hartnn):

do u know logarihms ?

hartnn (hartnn):

*logarithms

OpenStudy (emilyjones284):

no

hartnn (hartnn):

then which topic does this question belong to ?

OpenStudy (emilyjones284):

logarithms

OpenStudy (emilyjones284):

im just not sure how to do them because its review

hartnn (hartnn):

oh, okk.. first taking log on both sides, we get \(\log(4^{x+1})=\log 21\)

OpenStudy (emilyjones284):

and then what should i do?

hartnn (hartnn):

then we use the log property that \(\large \log a^b=b\log a\) so, what about \(\log 4^{x+1}\) ??

OpenStudy (emilyjones284):

ummm i dont understand what to do after that

hartnn (hartnn):

 got this ?